Calculate the control limits for a c chart with a process average equal to 4 nonconformities.
Answer #1

The 3-sigma control limits for c chart are given below:

UCL = c +3VC

CLc=c

LCL = -3Vc

where \bar{c} average number of process nonconformities and we are given \bar{c}=4.

Therefore,

UCL = 4+3V4

  = 4+3 x 2

  =446

=10

CLc=c

=4

LCL = -3Vc

  =4-3\sqrt{4}

=4-6

=-2=0

We took LCLc as we can't have a negative number of nonconformities for controlling.